An executive in an engineering firm earns a monthly salary plus a Christmas bonus of 6600 dollars. If she earned a total of 90300 dollars per year, what is her monthly salary in dollars.
Answer by Theo(13342)

she earned 90300 for the whole year.
this includes the christmas bonus of 6600.
subtract that from 90300 and you get a salary of 83700 for the year.
divide that by 12 and you get a monthly salary of 6975.
to calculate her yearly earnings, you would multiply the monthly salary by 12 and then add the christmas bonus on top of that.
12 * 6975 + 6600 = 90300.
put the commas where they belong and her monthly salary is 6,975 dollars.

algebraically, you would solve this as follows:

let x = her monthly salary.
her yearly salary is therefore 12*x.

your equation becomes 12*x + 6600 = 90300
subtract 6600 from both sides of the equation to get 12*x = 83700.
divide both sides of the equation by 12 to get x = 6975.